a square floor needs 400 square tiles each 10cm long. What is the area of the floor in square metres?

Answer by rwm(914)

400*10^2=cm^2
400*100=40000 cm^2
convert to square metres
200*200
it is 200 cm wide by 200 cm long
which is 2 metres by 2 metres=4 sq metres
another way
100*100 cm^2=1 square metre
40000/10000=4 square metres